# Geometric Neural Network based on Phase Space for BCI decoding

### v1

@@ Abstract
Objective: The integration of Deep Learning (DL) algorithms on brain signal analysis is still in its nascent stages compared to their success in fields like Computer Vision, especially in Brain-Computer Interface (BCI), where the brain activity is decoded to control external devices without requiring muscle control. Electroencephalography (EEG) is a widely adopted choice for designing BCI systems due to its non-invasive and cost-effective nature and excellent temporal resolution. Still, it comes at the expense of limited training data, poor signal-to-noise, and a large variability across and within-subject recordings. Finally, setting up a BCI system with many electrodes takes a long time, hindering the widespread adoption of reliable DL architectures in BCIs outside research laboratories. To improve adoption, we need to improve user comfort using, for instance, reliable algorithms that operate with few electrodes. Approach: Our research aims to develop a DL algorithm that delivers effective results with a limited number of electrodes. Taking advantage of the Augmented Covariance Method with SPDNet, we propose the SPDNet architecture and analyze its performance and computational impact, as well as the interpretability of the results. The evaluation is conducted on 5-fold cross-validation, using only three electrodes positioned above the Motor Cortex. The methodology was tested on nearly 100 subjects from several open-source datasets using the Mother Of All BCI Benchmark (MOABB) framework. Main results: The results of our SPDNet demonstrate that the augmented approach combined with the SPDNet significantly outperforms all the current state-of-the-art DL architecture in MI decoding. Significance: This new architecture is explainable, with a low number of trainable parameters and a reduced carbon footprint.
@@ 1 Introduction
This research focuses on developing a novel DL architecture, SPDNet, that outperforms the state-of-the-art classification when using a limited number of electrodes. Building on the Augmented Covariance Method (ACM) [7] that is an extension of the spatial covariance and the Symmetric Positive Definite (SPD) Neural Network -SPDNet [8], we study the SPDNet impact on the performance when using a reduced number of electrodes. We also conducted a comprehensive analysis of the model size, carbon emissions, and, ultimately, its explainability with respect to the standard SPDNet.
Spatial covariance, however, is not the only one that can be extracted from the EEG signal. Another potential candidate is coherence, which provides an alternative perspective on EEG signal characteristics. Historically, coherence has proven to be an inherently unstable feature, which is challenging to compute accurately, resulting in less robust results with respect to the spatial covariance. However, the use of information from both covariance and coherence has been shown to increase performance compared to methodologies relying solely on covariance [9]. We are thus interested to study coherence in combination with SPDNet and especially to use the ACM methodology adapted for the coherence matrix.
@@
The article is organized as follows: In section 2, we delve into the current state-of-the-art in Deep Learning (DL) EEG decoding, emphasizing distinctions from our approach. section 3 provides an overview of the theoretical foundations of our model and the considered datasets. The obtained results from the Within-session evaluation are presented in section 4. Subsequently, section 5 focuses on the method’s impact and current limitations, specifically focusing on its explainability and environmental impact. Finally, section 6 summarizes the findings of our study.
@@ 2.1 Machine Learning for EEG Decoding
Schirrmeister et al. [14] demonstrated that Deep Learning (DL) approaches, specifically ShallowNet and DeepNet can perform on par with conventional machine learning in decoding raw EEG data and can be trained end-to-end, eliminating several feature extraction steps. The application of DL algorithms enhances the generalization capability, enabling it to handle the inherent variability present in EEG signals effectively. This applicability of DL for EEG data is further corroborated by [18, 19, 20, 21, 15]. Despite their effectiveness, these DL approaches demand extensive parameter tuning, high power consumption, and large datasets for thorough evaluation. In contrast, our study proposes a streamlined approach, leveraging Riemannian geometry and functional connectivity derivatives, which are less demanding in terms of parameters, enhancing both efficiency and scalability.
@@ 2.2 Deep Riemannian Networks for EEG Decoding
Huang and Van Gool [22] work introduced SPDNet, a neural network that operates on SPD manifolds. This foundation was used by Ju et al. [35], who applied SPDNet in bio-signal classification, enhancing transfer learning. Building on this, Brooks et al. [36] adapted batch normalization for the SPD manifold, and Kobler et al. [25] further refined the batch normalization component for EEG Decoding. Pan et al. [24] added to this perspective by proposing attention components for the SPDNet zoo. Follow-up efforts included using residual layers [26, 31], filter bank inputs [23, 27, 30], mixing traditional convolution by channels [34], and constraining diffusion models [37, 38]. Contrary to our approach, these methods rely on adapting existing neural network components for the SPD manifold or preserving the dimensionality assumptions, while our method combines the dimensionality expansion of the phase space reconstruction on the SPD manifold.
@@ 2.3 Geometry Transformation in EEG data
Geometry Transformation (GT) in EEG analysis involves applying a transformation function to the data [6, 15, 39, 7]. This process commonly includes standard pre-processing transformations like resampling, band-passing, and filtering, which have been proven to enhance model performance and biomarker discovery [40, 41, 42, 43, 16]. Data augmentation, a regularization geometry transformation, has been acknowledged for its role in enhancing brain decoding in EEG, with its benefits varying depending on the task [44, 15]. While these transformations can improve the learning processing, these steps usually do not consider the non-stationary dynamics properties of the biological signal. Our approach broadens this scope by incorporating transformations that consider phase reconstruction in terms of its components of nonlinear dynamics.
@@ 3.1 EEG Decoding
We consider EEG signals as real-valued matrices. Let us denote each bio-signal window captured during a cognitive task on a given electrode by , a time serie of elements. We denote an epoch as , with and the corresponding cognitive task is denoted by . Here, represents the number of channels (electrodes), is the number of time points in the window, and indexes the trial , where is the total number of trials.
@@ 3.2 Phase space reconstruction for the EEG signal
Our method applies Takens’ theorem over the minimally pre-processed epoch-cropped time series, which allows for the understanding of a system’s multi-variable dynamics through a single observable in a -dimensional space by employing a phase embedding transformation as the initial function in our neural network [47, 48, 49]. This is made by constructing a phase space using a delay vector constructed from the original signal, thereby enabling the reconstruction of the system dynamics from these observations. Consider an embedding delay and a delay function , the delay vector is defined as:
where represents the embedding dimension, which dictates the order of magnitude of the phase space, and is the embedding delay, determining the temporal resolution of our analysis [50, 51]. As a result, embodies an embedding of the original phase space into a higher-dimensional space, enabling a detailed examination of the system’s dynamics.
Consider EEG window signals captured during a cognitive task as , and consider this signal produced by a nonlinear dynamical system. To create the higher dimensional space, we use the methodology proposed by Carrara and Papadopoulo [7], consisting of sliding measures over the observable with a subsequent concatenation. We define the lag function , and given the parameters and , is defined as:
This function is applied on each window signal .
@@
The MDOP algorithm emerged as a novel geometric for analyzing dynamical systems, as opposed to conventional statistical and information-theoretic methods, such as mutual information and continuity statistics. Central to MDOP is optimizing an embedding to ensure the reconstructed attractor is as expanded as possible while simultaneously reducing redundancy among the delay components. The algorithm employs a recursive approach, wherein each step of the embedding cycle identifies the lag that yields the highest beta statistics. This chosen lag is then utilized in the subsequent reconstruction phase. This iterative process continues until the algorithm obtains satisfying embedding dimensions , selected using the method of false nearest neighbors.
@@ 3.3 Riemannian Manifold
where represent the differential at point of the exponential function and similarly for . This formulation can be simplified if the reference matrix for the mapping is the identity matrix.
@@ 3.3.1 Symmetric Positive Definite Neural Networks Components
In a DL typical training process, optimization methods like Adam [60] are normally used for implementing the back-propagation procedure. However, in the context of SPDNet, we must the weights of of the BiMap layer in such a way that the new weights still belong to the Stiefel manifolds, i.e., they are still orthonormal matrices. To solve the problem, we use the RiemannAdam [61] optimization from the geoopt [62] library. In addition to the optimization strategy, we use the standard cross-entropy loss function as a loss function, a well-established and widely used objective function in classification tasks.
To summarize, the resulting architecture SPDNetuses enriched SPD matrices that encapsulate a broader spectrum of information than their traditional counterparts. In addition, such an approach allows, through the use of nonlinear systems theory, the reconstruction of a phase space that contains more information with respect to the one extracted from the original signal, allowing the use of fewer electrodes. The augmented SPD matrices require an SPDNet with a larger number of parameters, which is partly counteracted by using fewer electrodes. In particular, this representation accentuates how the input SPD matrix is modified from the initial dimension to .
@@
In the case of SPDNet, we set the subspace dimension of the BiMap layer at half of the input dimension [8]. However, for standard SPDNet, we maintained the subspace dimension at the same value as the original input dimension (although in this case the BiMap layer acts only as a rotation) in order to proceed with subsequent analyses concerning the explainability of the model. Anyway, diminishing the subspace dimension of half in the case of SPDNet consistently led to a reduction in performance. Note that the current implementation of SPDNet has a ReEig layer that is not scale-independent because of the parameter . To enhance the architecture’s scale independence, we have introduced a standardization procedure for the raw signal, bringing every channel to a zero mean and a unit standard deviation.
@@ 3.4 Datasets
We employed standard pre-processing steps designed for all datasets described in [10]. We applied band-pass filtering with the overlap-add method between Hz, artifact rejection, and electrode standardization, bringing every channel to a zero mean and a unit standard deviation. Additionally, we opted to use the complete epoch duration, although this duration varies among datasets, as shown in Table 1.
@@
Additionally, we focus on investigating our algorithm’s robustness with fewer electrode scenes. For this purpose, we selected a consistent set of three electrodes (C3, Cz, and C4) across datasets, all adhering to the 10-20 standard montage. Importantly, these electrodes are situated directly above the motor cortex, an area concentrating the most relevant information for MI-BCI classification [69].
@@ 3.5 Baseline Comparison
We compare the performance of our pipeline to to that of several state-of-the-arts DL Neural Networks used for BCI. All algorithms in our study were tested using the same set of three electrodes to maintain uniformity in our experimental approach. To standardize the models’ parameters, we initially resample the input time series to align the signal with the state of the art in order to use the parameter’s architecture of the original paper.
@@
EEGITNet [20], a neural network inspired by InceptionNet with parallel convolution layers with different scales. The parameters network has been designed to EEG signals at 128Hz.
@@
EEGNeX [21], a neural network inspired by EEGNet incorporation the key components from ConvNeXt. The parameters network has been designed to EEG signals at 128Hz.
@@ 4 Results
The first analysis is dedicated to contrasting the performance of SPDNet, with a specific focus on the exploitation of the covariance feature, against the state-of-the-art techniques in DL applied in EEG. The subsequent analysis, reported in the Appendix, shifts the focus towards the exploration of coherence as a feature for classification. Despite the ongoing investigations into coherence, we deliberately separated this analysis due to its comparatively lesser performance against the benchmarks achieved with covariance features. Nevertheless, the augmentation procedure is a significative methodology to improve the performance of both feature sets.
@@
For the first analysis, we tested the performance of SPDNet against the baseline comparison defined in 3.5 across a broad spectrum of datasets. The results are listed in Table 2, while the detailed statistical analysis can be found in Figure 2.
@@
Table 2 shows ROC AUC results for right-hand versus left-hand classification, where bold numbers represent the best score in each dataset, revealing a consistent trend wherein our algorithm SPDNet surpasses the DL state-of-the-art in the context of reduced datasets. This behavior is true for all datasets, with the notable exception of the Schirrmeister2017 dataset, where it is marginally surpassed by ShallowNet and DeepNet. It is important to note that the results of the latter two, however, are not statistically significant for this dataset. As the number of subjects per dataset is too small to reach decent statistical power for investigating the differences between pipelines, the results across datasets are aggregated to conduct a meta-analysis. Our SPDNet yields statistically better results than all other approaches as shown in Fig. 2 (b), where only significant interactions are displayed ( > 0.0.5), the reported values and color show the -values.
@@
The second relevant point is to see the impact of the augmentation procedure compared to the standard SPDNet. Figure 2 (a) elucidates this aspect by illustrating the relative performance enhancements of various models against the SPDNet standard. Notably, ShallowConvNet and DeepConvNet exhibit positive performance increments across the majority of the datasets examined. However, the augmentation procedure, as implemented by SPDNet, not only consistently yields positive relative improvement against the standard SPDNet across all datasets but also the improvement introduced by the augmentation procedure turns out to be so significant that it even outperforms methods that initially surpassed the SPDNet benchmark. This improvement is not only demonstrated by the relative improvement with respect to the SPDNet standard but also is evident through the meta-analysis comparison(refer to Fig 2 (c), (d) and (e)).
@@
Ultimately, the statistical analysis conducted using Figure 2 (b) allows us to demonstrate that SPDNet outperforms the state-of-the-art DL methodology in a reduced dataset context. Its versatility and reliability, particularly in data-limited scenarios, position the method as a promising solution for real-world applications.
@@
| DeepNet [14] | 75.80 15.45 | 72.80 19.48 | 63.13 14.45 | 73.04 15.67 | 73.97 18.07 | 91.74 7.00 |
| ShallowNet [14] | 75.85 15.41 | 72.17 18.61 | 64.14 13.03 | 73.59 15.19 | 75.36 15.69 | 88.03 8.55 |
| EEGNet [18] | 70.64 19.87 | 70.27 18.91 | 60.23 14.98 | 69.80 16.51 | 71.94 17.80 | 88.95 7.84 |
| EEGTCNet [19] | 65.98 17.25 | 66.86 18.41 | 56.43 12.31 | 67.87 17.36 | 65.94 15.69 | 81.47 11.66 |
| EEGITNet [20] | 66.64 13.98 | 64.93 14.49 | 54.68 11.97 | 62.98 16.19 | 56.97 17.66 | 72.82 12.78 |
| EEGNeX [21] | 68.86 17.27 | 68.29 17.85 | 56.64 12.83 | 64.26 17.58 | 58.73 19.40 | 80.25 15.55 |
| SPDNet [8] | 71.02 15.64 | 70.15 16.85 | 59.95 12.61 | 67.40 13.01 | 67.04 17.62 | 88.85 8.05 |
| SPDNet (Our) | 75.98 17.00 | 80.46 16.64 | 66.00 13.29 | 72.02 13.07 | 78.01 19.64 | 94.92 3.34 |

### v2

# Geometric Neural Network based on Phase Space for BCI-EEG decoding

@@ Abstract
Objective: The integration of Deep Learning (DL) algorithms on brain signal analysis is still in its nascent stages compared to their success in fields like Computer Vision, especially in Brain-Computer Interface (BCI), where the brain activity is decoded to control external devices without requiring muscle control. Electroencephalography (EEG) is a widely adopted choice for designing BCI systems due to its non-invasive and cost-effective nature and excellent temporal resolution. Still, it comes at the expense of limited training data, poor signal-to-noise, and a large variability across and within-subject recordings. Finally, setting up a BCI system with many electrodes takes a long time, hindering the widespread adoption of reliable DL architectures in BCIs outside research laboratories. To improve adoption, we need to improve user comfort using, for instance, reliable algorithms that operate with few electrodes. Approach: Our research aims to develop a DL algorithm that delivers effective results with a limited number of electrodes. Taking advantage of the Augmented Covariance Method with SPDNet, we propose the Phase-SPDNet architecture and analyze its performance and computational impact, as well as the interpretability of the results. The evaluation is conducted on 5-fold cross-validation, using only three electrodes positioned above the Motor Cortex. The methodology was tested on nearly 100 subjects from several open-source datasets using the Mother Of All BCI Benchmark (MOABB) framework. Main results: The results of our Phase-SPDNet demonstrate that the augmented approach combined with the SPDNet significantly outperforms all the current state-of-the-art DL architecture in MI decoding. Significance: This new architecture is explainable, with a low number of trainable parameters and a reduced carbon footprint.
@@ 1 Introduction
This research focuses on developing a novel DL architecture, Phase-SPDNet , that outperforms the state-of-the-art classification when using a limited number of electrodes. Building on the Augmented Covariance Method (ACM) [7] that is an extension of the spatial covariance and the Symmetric Positive Definite (SPD) Neural Network -SPDNet [8], we study the Phase-SPDNet impact on the performance when using a reduced number of electrodes. We also conducted a comprehensive analysis of the model size, carbon emissions, and, ultimately, its explainability with respect to the standard SPDNet.
Spatial covariance, however, is not the only one that can be extracted from the EEG signal. Another potential candidate is coherence, which provides an alternative perspective on EEG signal characteristics. Historically, coherence has proven to be an inherently unstable feature, which is challenging to compute accurately, resulting in less robust results with respect to the spatial covariance. However, the use of information from both covariance and coherence has been shown to increase performance as we are considering an estimator of the interactions between brain areas/electrodes [9]. We are thus interested to study coherence in combination with SPDNet and especially to use the ACM methodology adapted for the coherence matrix.
@@
The article is organized as follows: In section 2, we examine the current state-of-the-art in Deep Learning (DL) EEG decoding, emphasizing distinctions from our approach. section 3 provides an overview of the theoretical foundations of our model and the considered datasets. The obtained results from the Within-session evaluation are presented in section 4. Subsequently, section 5 focuses on the method’s impact and current limitations, specifically focusing on its explainability and environmental impact. Finally, section 6 summarizes the findings of our study.
@@ 2.1 Machine Learning for EEG Decoding
Schirrmeister et al. [14] demonstrated that Deep Learning (DL) approaches, specifically ShallowNet and DeepNet can perform on par with conventional machine learning in decoding raw EEG data and can be trained end-to-end, eliminating several feature extraction steps. The application of DL algorithms enhances the generalization capability, enabling it to handle the inherent variability present in EEG signals effectively. This applicability of DL for EEG data is further corroborated by [18, 19, 20, 21, 15]. Despite their effectiveness, these DL approaches demand extensive parameter tuning, high power consumption, and large datasets for thorough evaluation. In contrast, our study proposes a streamlined approach, leveraging Riemannian geometry and functional connectivity derivatives, which are less demanding in terms of parameters, enhancing both efficiency and scalability. The efficiency and scalability stems from the geometric transformation these methods apply, which reduces the temporal signal’s dimensionality from channel and time dimensions to channel-by-channel dimensions.
@@ 2.2 Deep Riemannian Networks for EEG Decoding
Huang and Van Gool [22] work introduced SPDNet, a neural network that operates on SPD manifolds. This foundation was used by Ju et al. [35], who applied SPDNet in bio-signal classification, enhancing transfer learning. Building on this, Brooks et al. [36] adapted batch normalization for the SPD manifold, and Kobler et al. [25] further refined the batch normalization component for EEG Decoding. Pan et al. [24] added to this perspective by proposing attention components for the category of the SPDNet neural networks. Follow-up efforts included using residual layers [26, 31], filter bank inputs [23, 27, 30], mixing traditional convolution by channels [34], and constraining diffusion models [37, 38]. Contrary to our approach, these methods rely on adapting existing neural network components for the SPD manifold or preserving the dimensionality assumptions, while our method combines the dimensionality expansion of the phase space reconstruction on the SPD manifold.
@@ 2.3 Geometry Transformation in EEG data
Geometry Transformation (GT) [6, 15, 39, 7] in EEG analysis involves standard pre-processing transformations like resampling, band-passing, and filtering, which have been proven to enhance model performance and biomarker discovery [40, 41, 42, 43, 16]. Data augmentation, a regularization geometry transformation, has been acknowledged for its role in enhancing brain decoding in EEG, with its benefits varying depending on the task [44, 15]. While these transformations can improve the learning processing, these steps usually do not consider the non-stationary dynamics properties of the biological signal. Our approach broadens this scope by incorporating transformations that consider phase reconstruction in terms of its components of nonlinear dynamics.
@@
2.4 Functional Connectivity for EEG Decoding
Beyond the framework of processing the raw brain signal data, researchers in the neuroscience community have been interested in identifying the relationship between brain areas using Functional Connectivity (FC) features [47, 9]. These features demonstrate that electrode interactions are effective for BCI tasks [9, 47, 48, 49, 50, 51, 52, 53, 54]. Most EEG decoding studies using FC employ brain interaction estimators to construct adjacency matrices for each trial, such as covariance, correlation, and coherence, often combined with dimension reduction techniques. We refer to [47] for an extensive discussion on FC estimators for brain signals.
@@ 3.1 EEG Decoding
We consider EEG signals as real-valued matrices. Let us denote each bio-signal window captured during a cognitive task (called an epoch) by . Here, represents the number of channels (electrodes), is the number of time points in the window, and indexes the trial , where is the total number of trials. Each epoch corresponds to a specific cognitive task denoted by .
@@ 3.2 Phase space reconstruction for the EEG signal
Our method applies Takens’ theorem over the minimally pre-processed epoch-cropped time series, which allows for the understanding of a system’s multi-variable dynamics through a single observable in a -dimensional space by employing a phase embedding transformation as the initial function in our neural network [55, 56, 57]. This is made by constructing a phase space using a delay vector constructed from the original signal, thereby enabling the reconstruction of the system dynamics from these observations. Following the approach proposed by Carrara and Papadopoulo [7], it is possible to define a delay function for . This function allows the concatenation of measures on a sliding window over the observable in a delay vector defined as:
where represents the embedding dimension, which dictates the order of magnitude of the phase space, and is the embedding delay, determining the temporal resolution of our analysis [58, 59]. As a result, embodies an embedding of the original phase space into a higher-dimensional space, enabling a detailed examination of the system’s dynamics. This formulation considers the EEG signal as produced by a nonlinear dynamical system. Concerning the dimension, if the original epoch have dimension the matrix . This function is applied on each window signal . An illustration of the process, with and is present in Figure 1
@@
The MDOP algorithm emerged as a novel geometric method for analyzing dynamical systems, as opposed to conventional statistical and information-theoretic methods, such as mutual information and continuity statistics. Central to MDOP is optimizing an embedding to ensure the reconstructed attractor is as expanded as possible while simultaneously reducing redundancy among the delay components. The algorithm employs a recursive approach, wherein each step of the embedding cycle identifies the lag that yields the highest beta statistics. This chosen lag is then utilized in the subsequent reconstruction phase. This iterative process continues until the algorithm obtains satisfying embedding dimensions , selected using the method of false nearest neighbors.
@@ 3.3 Riemannian Manifold
where represent the differential at point of the exponential function and similarly for . This formulation can be simplified if the reference matrix for the mapping is the identity matrix. Equation 4 presents two quantities, the exponential and logarithmic maps, which are used to transition from Euclidean space to the Riemannian surface.
@@ 3.3.1 Symmetric Positive Definite Neural Networks Components
In a DL typical training process, optimization methods like Adam [68] are normally used for implementing the back-propagation procedure. However, in the context of SPDNet, we must learn the weights of of the BiMap layer in such a way that the new weights still belong to the Stiefel manifolds, i.e., they are still orthonormal matrices. To solve the problem, we use the RiemannAdam [69] optimization from the geoopt [70] library. In addition to the optimization strategy, we use the standard cross-entropy loss function as a loss function, a well-established and widely used objective function in classification tasks.
To summarize, the resulting architecture Phase-SPDNet uses enriched SPD matrices that encapsulate a broader spectrum of information than their traditional counterparts thanks to the additional layer that implements the phase space reconstruction. In addition, such an approach allows, through the use of nonlinear systems theory, the reconstruction of a phase space that contains more information with respect to the one extracted from the original signal, allowing the use of fewer electrodes. The augmented SPD matrices require an SPDNet with a larger number of parameters, which is partly counteracted by using fewer electrodes. In particular, this representation accentuates how the input SPD matrix is modified from the initial dimension to .
@@
In the case of Phase-SPDNet , we set the subspace dimension of the BiMap layer at half of the input dimension [8]. However, for standard SPDNet, we maintained the subspace dimension at the same value as the original input dimension (although in this case the BiMap layer acts only as a rotation) in order to proceed with subsequent analyses concerning the explainability of the model. Anyway, diminishing the subspace dimension of half in the case of SPDNet consistently led to a reduction in performance. Note that the current implementation of SPDNet has a ReEig layer that is not scale-independent because of the parameter . To enhance the architecture’s scale independence, we have introduced a standardization procedure for the raw signal, bringing every channel to a zero mean and a unit standard deviation.
In short, Phase SPDNet processes the SPD matrix that contains the spatial information and the temporal evolution of the signal, while SPDNet handles only the spatial covariance.
@@ 3.4 Datasets
We employed standard pre-processing steps designed for all datasets described in [10]. We applied band-pass filtering with the overlap-add method between Hz [77, 78], artifact rejection, and electrode standardization, bringing every channel to a zero mean and a unit standard deviation. Additionally, we opted to use the complete epoch duration, although this duration varies among datasets, as shown in Table 1. The pre-processing steps just mentioned are the same for all the pipelines considered in our study.
@@
Additionally, we focus on investigating our algorithm’s robustness with fewer electrode settings. For this purpose, our channel selection process was guided by two constraints: ensuring neurophysiological relevance and maintaining consistent electrode selection across datasets. We selected electrodes positioned over the sensorimotor region, particularly in the central (C) and centro-parietal (CP) lines, which are actively involved in the tasks. This choice is strongly supported by Brain-Computer Interface (BCI) research [78, 79, 80, 81, 82, 83, 84]. For consistency, we used EEG montages with varying numbers of electrodes, from 3 to 128. We standardized our analysis by using the minimal overlapping electrodes common to all montages: C3, Cz, and C4.
@@ 3.5 Baseline Comparison
We compare the performance of our pipeline to that of several state-of-the-art DL Neural Networks used for BCI. All algorithms in our study were tested using the same set of three electrodes to maintain uniformity in our experimental approach. To standardize the models’ parameters, we initially resample the input time series to align the signal with the state of the art in order to use the parameter’s architecture of the original paper.
@@
EEGITNet [20], a neural network inspired by InceptionNet with parallel convolution layers with different scales. The network has been designed for EEG signals sampled at 128Hz.
@@
EEGNeX [21], a neural network inspired by EEGNet incorporation the key components from ConvNeXt. The network has been designed for EEG signals sampled at 128Hz.
7.
DynSpat-EEGNet [85], a neural network based on Dynamic spatial filtering with the EEGNet as head. The network has been designed for EEG signals sampled at 250Hz.
8.
DynSpat-ShallowNet [85], a neural network based on Dynamic spatial filtering with the EEGNet as head. The network has been designed for EEG signals sampled at 250Hz.
@@ 4 Results
The first analysis is dedicated to contrasting the performance of Phase-SPDNet , with a specific focus on the exploitation of the covariance feature, against the state-of-the-art techniques in DL applied in EEG. The subsequent analysis, reported in the Appendix, shifts the focus towards the exploration of coherence as a feature for classification. Despite the ongoing investigations into coherence, we deliberately separated this analysis due to its comparatively lesser performance against the benchmarks achieved with covariance features. Nevertheless, the augmentation procedure is a significative methodology to improve the performance of both feature sets.
@@
For the first analysis, we tested the performance of Phase-SPDNet against the baseline comparison defined in 3.5 across a broad spectrum of datasets. The results are listed in Table 2, while the detailed statistical analysis can be found in Figure 3.
@@
Table 2 shows ROC AUC results for right-hand versus left-hand classification, where bold numbers represent the best score in each dataset, revealing a consistent trend wherein our algorithm Phase-SPDNet surpasses the DL state-of-the-art in the context of reduced datasets. This behavior is true for all datasets, with the notable exception of the Schirrmeister2017 dataset, where it is marginally surpassed by ShallowNet, DeepNet and DynSpat-ShallowNet. It is important to note that the results of the first two, however, are not statistically significant for this dataset, as presented in Figure 3 (c), (d). The results obtained in this dataset by DynSpat-ShallowNet is statistically superior, as shown in Figure 3 (e). However, we are interested in computing a statistic behavior considering all datasets under consideration. As the number of subjects per dataset is actually too small to reach a decent statistical power for investigating the differences between pipelines, the results across datasets are aggregated to conduct a meta-analysis. Our Phase-SPDNet yields statistically better results than all other approaches as shown in Fig. 3 (b), where only significant interactions are displayed ( > 0.0.5), the reported values and color show the -values.
@@
The second relevant point is to see the impact of the augmentation procedure compared to the standard SPDNet. Figure 3 (a) elucidates this aspect by illustrating the relative performance enhancements of various models against the SPDNet standard. Notably, ShallowNet and DeepNet exhibit positive performance increments across the majority of the datasets examined. However, the augmentation procedure, as implemented by Phase-SPDNet , not only consistently yields positive relative improvement against the standard SPDNet across all datasets but also the improvement introduced by the augmentation procedure turns out to be so significant that it even outperforms methods that initially surpassed the SPDNet benchmark. This improvement is not only demonstrated by the relative improvement with respect to the SPDNet standard but also is evident through the meta-analysis comparison (refer to Fig 3 (c), (d) and (e)).
@@
Ultimately, the statistical analysis conducted using Figure 3 (b) allows us to demonstrate that Phase-SPDNet outperforms the state-of-the-art DL methodology in a reduced dataset context. Its versatility and reliability, particularly in data-limited scenarios, position the method as a promising solution for real-world applications.
@@
| DeepNet [14] (2017) | 75.80 15.45 | 72.80 19.48 | 63.13 14.45 | 73.04 15.67 | 73.97 18.07 | 91.74 7.00 |
| ShallowNet [14] (2017) | 75.85 15.41 | 72.17 18.61 | 64.14 13.03 | 73.59 15.19 | 75.36 15.69 | 88.03 8.55 |
| EEGNet [18] (2018) | 70.64 19.87 | 70.27 18.91 | 60.23 14.98 | 69.80 16.51 | 71.94 17.80 | 88.95 7.84 |
| EEGTCNet [19] (2020) | 65.98 17.25 | 66.86 18.41 | 56.43 12.31 | 67.87 17.36 | 65.94 15.69 | 81.47 11.66 |
| EEGITNet [20] (2022) | 66.64 13.98 | 64.93 14.49 | 54.68 11.97 | 62.98 16.19 | 56.97 17.66 | 72.82 12.78 |
| EEGNeX [21] (2022) | 68.86 17.27 | 68.29 17.85 | 56.64 12.83 | 64.26 17.58 | 58.73 19.40 | 80.25 15.55 |
| DynSpat+EEGNet [85] (2022) | 68.06 18.58 | 76.61 16.81 | 61.51 13.74 | 70.62 16.53 | 72.88 19.33 | 87.85 7.01 |
| DynSpat+ShallowNet [85] (2022) | 70.95 18.78 | 78.95 17.23 | 65.52 14.65 | 75.07 14.40 | 74.70 19.12 | 87.33 8.87 |
| SPDNet [8] (2017) | 71.02 15.64 | 70.15 16.85 | 59.95 12.61 | 67.40 13.01 | 67.04 17.62 | 88.85 8.05 |
| Phase-SPDNet (Our) | 75.98 17.00 | 80.46 16.64 | 66.00 13.29 | 72.02 13.07 | 78.01 19.64 | 94.92 3.34 |